# No Dimensional Sampling Coresets for Classification

### v1

@@ 2.1 Kernelization of the Results
$\mathcal{L}^{H}_{\phi,k}=\left\{\ell^{H}_{\phi,k}(w,\cdot)=\phi(\boldsymbol{K}(w,\cdot))+\frac{1}{k}\boldsymbol{K}(w,w)\colon w\in\mathcal{X}\right\}$ (8)
@@
¯$\bar{\mathcal{L}}^{H}_{\phi,k}=\left\{\bar{\ell}^{H}_{\phi,k}(w,\cdot)=\phi(-\boldsymbol{K}(w,\cdot))+\frac{1}{k}\boldsymbol{K}(w,w)\colon w\in\mathcal{X}\right\}$ (9)
To keep things simpler, we occasionally employ $\|X\|_{H}$ and $\langle\cdot,\cdot\rangle_{H}$ instead of $\sqrt{\boldsymbol{K}(x,x)}$ and $\boldsymbol{K}(\cdot,\cdot)$ respectively.
Remark 9
@@
Theorem 10 (Theorem 5, Kernelized Representation)
For a well-behaved probability measure $P$ with respect to a Hilbert space $H$ of real-valued functions defined on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, if $s(\cdot):\mathcal{X}\longrightarrow(1,\infty)$ is an upper sensitivity function for $\mathcal{L}^{H}_{\phi,k}$ with total sensitivity $S$ and $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$, then any $s$-sensitivity sample $x_{1},\ldots,x_{m}$ from $\mathcal{X}$ with weights $u_{i}=\frac{S}{ms(x_{i})}$ provides an $\varepsilon$-coreset for $(\mathcal{X},\ P,\ \mathcal{X},\ \ell^{H}_{\phi,k})$ with probability at least $1-\delta$, where
@@
Considering $\mathbb{R}^{d}$ as a Hilbert space equipped with linear kernel, we derive Theorem 5 as a corollary of this theorem. In the following, we present various applications derived from this theorem.
Theorem 11
Let us consider $H$, a reproducing kernel Hilbert space of real-valued functions defined on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, and $P$, a probability measure over $\mathcal{X}$. Given the conditions outlined in the initial $4$ columns of Table 3, any $s$-sensitivity sample of size $m$ (sampling according to the distribution provided in Column 5) guarantees an $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{X},\mathcal{L}^{H}_{\phi,k})$ with a probability of at least $1-\delta$.
@@
$\sigma(\pm\cdot)$ $\operatorname*{\mathbb{E}}\limits_{x\sim P}\|x\|_{H}^{2}\leq E^{2}_{1}$ $O(E_{1}^{2}k)$ $O(E_{1}^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Theorem 24
@@
$\operatorname{{\rm logistic}}(\pm\cdot)$ $P\left(\|x\|_{H}\geq A\right)=0$ $O(\frac{A^{2}k}{\sqrt{\log(A^{2}k)}})$ $O(A^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Theorem 28
@@
$\operatorname{{\rm logistic}}(\pm\cdot)$ $\operatorname*{\mathbb{E}}\limits_{x\sim P}\|x\|_{H}^{2}\leq E_{1}$ $O(\frac{E_{1}^{2}k}{\sqrt{\log(E_{1}^{2}k)}})$ $O(E_{1}^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Theorem 31
$\operatorname{{\rm svm}}(\pm\cdot)$ $P\left(\|x\|_{H}\geq A\right)=0$ $O(A^{2}k)$ $O(A^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Theorem 36
@@
$\operatorname{{\rm svm}}(\pm\cdot)$ $\operatorname*{\mathbb{E}}\limits_{x\sim P}\|x\|_{H}^{2}\leq E_{1}$ $O(E_{1}^{2}k)$ $O(E_{1}^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Theorem 40
2.1.1 Kernel Density Estimate
@@
Let $P$ be a probability measure over $\mathbb{R}^{d}$ and $\boldsymbol{K}:\mathbb{R}^{d}\times\mathbb{R}^{d}\longrightarrow\mathbb{R}$ be a kernel, for instance the Gaussian kernel $\boldsymbol{K}(x,w)=\exp(-\|x-w\|^{2})$. At any point $w\in\mathbb{R}^{d}$, the kernel density estimate is defined ad $\text{KDE}_{P}(w)=\int_{x\in\mathbb{R}^{d}}\boldsymbol{K}(x,w){\rm d}P(x).$ When we have finite number of points $X=\{x_{1},\ldots,x_{n}\}$ given as the data set, we can assume that $P$ is a uniform probability measure over $X$, i.e., $P(x=x_{i})=\frac{1}{n}$. In this scenario, we use $\text{KDE}_{X}$ instead of $\text{KDE}_{P}$ and and the expression for $\text{KDE}_{X}$ is given by
$\text{KDE}_{X}(w)=\frac{1}{n}\sum_{x\in X}\boldsymbol{K}(x,w).$
@@
The evaluation of $\text{KDE}_{X}$ demands O(n) time, which can become impractical for massive datasets. Therefore, a frequently employed approach is to substitute $X$ with a significantly smaller dataset $Y$, allowing $\text{KDE}_{Y}$ to function as an approximation for $\text{KDE}_{X}$. Formally, for a given $\varepsilon\in(0,1)$, we look for a small size $Y$ such that
$\sup_{w\in\mathbb{R}^{d}}|\text{KDE}_{X}(w)-\text{KDE}_{Y}(w)|\leq\varepsilon.$
@@
This challenge has been thoroughly explored and investigated in various studies such as (Phillips and Tai, 2020; Tai, 2022; Charikar et al., 2024; Bobrowski et al., 2017; Taylor, 2018; Gretton et al., 2012; Phillips et al., 2015; Rinaldo and Wasserman, 2010; Scott, 2015; Silverman, 1986; Zheng and Phillips, 2015). Our primary result in this context is presented in the following theorem, restated as Corollary 37 along with its proof.
Theorem 12
@@
Let $\boldsymbol{K}:\mathbb{R}^{d}\times\mathbb{R}^{d}\longrightarrow(0,1]$ be a reproducing kernel, i.e., a kernel associated with an RKHS, and $P$ be a probability measure over $\mathbb{R}^{d}$. For $\varepsilon,\delta\in(0,1)$, there exists a universal constant $C$ (independent of $d$ and $\boldsymbol{K}$) such that if $m\geq\frac{C}{\varepsilon^{2}}\log\frac{1}{\delta}$, then, with probability at least $1-\delta$, for any iid random sample $X=\{x_{1},\ldots,x_{m}\}$ based on $P$, we have
[…]
@@
Note that here, we treat $d$ as a variable, and importantly, our bound is independent of $d$. The Gaussian, Laplacian, and Exponential kernels are typically chosen for study. Compared to the state-of-the-art results, we can reproduce the sampling result of $O(1/\varepsilon^{2})$ provided in the appendix of (Lopez-Paz et al., 2015) (stated with constant probability of failure $\delta$), which is the same bound as the greedy selection method (Lacoste-Julien et al., 2015), and is optimal if we do not bound the dimension (Phillips and Tai, 2018a). Nevertheless, there are some improvements in these results, including when $d<1/\varepsilon^{2}$, such as works by (Phillips and Tai, 2018a, 2020; Karnin and Liberty, 2019), and more recent papers by (Tai, 2022; Charikar et al., 2024).
3 Main Tools
@@
3.1 Proof of Theorem 2
We begin by revisiting McDiarmid’s inequality, which serves as a concentration inequality that bounds the deviation of a function of several random variables from its expected value, under the assumption that changing any one of the variables can only have a limited effect on the function. If $(\mathcal{X}_{1},\Sigma_{1},P_{1}),\ldots,(\mathcal{X}_{n},\Sigma_{1},P_{n})$ are probability measure spaces, then $P=\prod_{i=1}^{n}P_{i}$ is a probability measure over $\prod_{i=1}^{n}\mathcal{X}_{i}$. A function $f:\prod_{i=1}^{n}\mathcal{X}_{i}\longrightarrow\mathbb{R}$ satisfies the bounded differences property if there are constants $c_{1},c_{2},\dots,c_{n}$ such that for all $i\in[n]$, and for all $x_{1}\in{\mathcal{X}}_{1},\,x_{2}\in{\mathcal{X}}_{2},\,\ldots,\,x_{n}\in{\mathcal{X}}_{n}$
@@ 3.1 Proof of Theorem 2
[…]
@@
Theorem 14 (Theorem 2, Restated)
@@
Proof  In view of Equation (3), we need to show that, for any iid sample $x_{1},\ldots,x_{m}\in\mathcal{X}$ according to the sensitivity normalized distribution $q(x)=\frac{s(x)}{S}p(x)$, with probability at least $1-2\exp\left(-{\frac{2mt^{2}}{S^{2}}}\right)$,
@@
(*(*$\qquad$( * ) $=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{x^{\prime}_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{f\in\mathcal{F}}\left[\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\left(T_{f}(x^{\prime}_{i})-T_{f}(x_{i})\right)\right]$
@@
The equality marked by (*)(*)( * ) is true since $\sigma_{i}$ indeed exchanges $x_{i}$ and $x_{i}^{\prime}$ together, and since they are iid samples from $q$ it cannot change the expectation. With the same approach, we would have
@@
which concludes $\max\{A,B\}\leq 2\mathcal{R}^{q}_{m}(\mathcal{T}_{\mathcal{F}})$ completing the proof.
3.2 Contraction Lemma
When dealing with Rademacher complexity, we can leverage the helpful lemma known as Talagrand’s Contraction Lemma (refer to Theorem 4.12 in (Ledoux and Talagrand, 1991)). This lemma, in particular, establishes that $\mathcal{R}_{m}(\phi\circ\mathcal{F})\leq L\mathcal{R}_{m}(\mathcal{F})$, where $\phi\colon\mathbb{R}\longrightarrow\mathbb{R}$ is an $L$-Lipschitz function, and $\phi\circ\mathcal{F}=\{\phi\circ f\colon f\in\mathcal{F}\}$. Our requirement extends beyond this lemma to a generalization.
Lemma 15

### v2

@@ A.1 Kernelization of the Results
$\mathcal{L}^{H}_{\phi,k}=\left\{\ell^{H}_{\phi,k}(w,\cdot)=\phi(\mathsf{K}(w,\cdot))+\frac{1}{k}\mathsf{K}(w,w)\colon w\in\mathcal{X}\right\}$ (13)
@@
¯$\bar{\mathcal{L}}^{H}_{\phi,k}=\left\{\bar{\ell}^{H}_{\phi,k}(w,\cdot)=\phi(-\mathsf{K}(w,\cdot))+\frac{1}{k}\mathsf{K}(w,w)\colon w\in\mathcal{X}\right\}$ (14)
To keep things simpler, we occasionally employ $\|X\|_{H}$ and $\langle\cdot,\cdot\rangle_{H}$ instead of $\sqrt{\mathsf{K}(x,x)}$ and $\mathsf{K}(\cdot,\cdot)$ respectively.
Remark A.2.
@@
Theorem A.3 (Theorem 1.4, Kernelized Representation).
For a well-behaved probability measure $P$ with respect to a Hilbert space $H$ of real-valued functions defined on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, if $s(\cdot):\mathcal{X}\longrightarrow(1,\infty)$ is an upper sensitivity function for $\mathcal{L}^{H}_{\phi,k}$ with total sensitivity $S$ and $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$, then any $s$-sensitivity sample $x_{1},\ldots,x_{m}$ from $\mathcal{X}$ with weights $u_{i}=\frac{S}{ms(x_{i})}$ provides an $\varepsilon$-coreset for $(\mathcal{X},\ P,\ \mathcal{X},\ \ell^{H}_{\phi,k})$ with probability at least $1-\delta$, where
@@
Considering $\mathbb{R}^{d}$ as a Hilbert space equipped with linear kernel, we derive Theorem 1.4 as a corollary of this theorem. In the following, we present various applications derived from this theorem.
Theorem A.4.
Let us consider $H$, a reproducing kernel Hilbert space of real-valued functions defined on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, and $P$, a probability measure over $\mathcal{X}$. Given the conditions outlined in the initial $4$ columns of Table 3, any $s$-sensitivity sample of size $m$ (sampling according to the distribution provided in Column 5) guarantees an $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{X},\mathcal{L}^{H}_{\phi,k})$ with a probability of at least $1-\delta$.
@@
$\sigma(\pm\cdot)$ $\operatorname*{\mathbb{E}}\limits_{x\sim P}\|x\|_{H}^{2}\leq E^{2}_{1}$ $O(E_{1}^{2}k)$ $O(E_{1}^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Thm D.6
@@
$\operatorname{{\rm logistic}}(\pm\cdot)$ $P\left(\|x\|_{H}\geq A\right)=0$ $O(\frac{A^{2}k}{\sqrt{\log(A^{2}k)}})$ $O(A^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Thm D.10
@@
$\operatorname{{\rm logistic}}(\pm\cdot)$ $\operatorname*{\mathbb{E}}\limits_{x\sim P}\|x\|_{H}^{2}\leq E_{1}$ $O(\frac{E_{1}^{2}k}{\sqrt{\log(E_{1}^{2}k)}})$ $O(E_{1}^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Thm D.13
$\operatorname{{\rm svm}}(\pm\cdot)$ $P\left(\|x\|_{H}\geq A\right)=0$ $O(A^{2}k)$ $O(A^{2}k)$ $p$ $\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$ Thm D.18
@@
$\operatorname{{\rm svm}}(\pm\cdot)$ $\operatorname*{\mathbb{E}}\limits_{x\sim P}\|x\|_{H}^{2}\leq E_{1}$ $O(E_{1}^{2}k)$ $O(E_{1}^{2}k)$ $\frac{s}{S}p$ $\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta})\right)$ Thm D.22
A.1.1 Kernel Density Estimate
@@
Let $P$ be a probability measure over $\mathbb{R}^{d}$ and $\mathsf{K}:\mathbb{R}^{d}\times\mathbb{R}^{d}\longrightarrow\mathbb{R}$ be a kernel, for instance the Gaussian kernel $\mathsf{K}(x,w)=\exp(-\|x-w\|^{2})$. At any point $w\in\mathbb{R}^{d}$, the kernel density estimate is defined ad ${\rm KDE}_{P}(w)=\int_{x\in\mathbb{R}^{d}}\mathsf{K}(x,w){\rm d}P(x).$ When we have finite number of points $X=\{x_{1},\ldots,x_{n}\}$ given as the data set, we can assume that $P$ is a uniform probability measure over $X$, i.e., $P(x=x_{i})=\frac{1}{n}$. In this scenario, we use ${\rm KDE}_{X}$ instead of ${\rm KDE}_{P}$ and and the expression for ${\rm KDE}_{X}$ is given by
${\rm KDE}_{X}(w)=\frac{1}{n}\sum_{x\in X}\mathsf{K}(x,w).$
@@
The evaluation of ${\rm KDE}_{X}$ demands O(n) time, which can become impractical for massive datasets. Therefore, a frequently employed approach is to substitute $X$ with a significantly smaller dataset $Y$, allowing ${\rm KDE}_{Y}$ to function as an approximation for ${\rm KDE}_{X}$. Formally, for a given $\varepsilon\in(0,1)$, we look for a small size $Y$ such that
$\sup_{w\in\mathbb{R}^{d}}|{\rm KDE}_{X}(w)-{\rm KDE}_{Y}(w)|\leq\varepsilon.$
@@
This challenge has been thoroughly explored and investigated in various studies such as (Phillips & Tai, 2020; Tai, 2022; Charikar et al., 2024; Bobrowski et al., 2017; Taylor, 2018; Gretton et al., 2012; Phillips et al., 2015; Rinaldo & Wasserman, 2010; Scott, 2015; Silverman, 1986; Zheng & Phillips, 2015). Our primary result in this context is presented in the following theorem, restated as Corollary D.19 along with its proof.
Theorem A.5.
@@
Let $\mathsf{K}:\mathbb{R}^{d}\times\mathbb{R}^{d}\longrightarrow(0,1]$ be a reproducing kernel, i.e., a kernel associated with an RKHS, and $P$ be a probability measure over $\mathbb{R}^{d}$. For $\varepsilon,\delta\in(0,1)$, there exists a universal constant $C$ (independent of $d$ and $\mathsf{K}$) such that if $m\geq\frac{C}{\varepsilon^{2}}\log\frac{1}{\delta}$, then, with probability at least $1-\delta$, for any iid random sample $X=\{x_{1},\ldots,x_{m}\}$ based on $P$, we have
$\sup_{w\in\mathbb{R}^{d}}|{\rm KDE}_{P}(w)-{\rm KDE}_{X}(w)|\leq\varepsilon.$
@@
Note that here, we treat $d$ as a variable, and importantly, our bound is independent of $d$. The Gaussian, Laplacian, and Exponential kernels are typically chosen for study. Compared to the state-of-the-art results, we can reproduce the findings of (Lopez-Paz et al., 2015) and (Lacoste-Julien et al., 2015). Nevertheless, there are some improvements in these results, such as works by (Phillips & Tai, 2018a, 2020; Karnin & Liberty, 2019), and more recent papers by (Tai, 2022; Charikar et al., 2024). It is worth noting that some of these results consider $d$ as a constant, while others treat it as a variable.
Appendix B Related Works
In machine learning, a common scenario involves having a set of $n$ data points and the goal of minimizing a data-driven objective loss function. At times, for scalability reasons, it becomes necessary to choose a smaller subset of $m\ll n$ points. The aim is to minimize the objective function on these points, potentially using non-uniform weights for the selected points. This approach aims to achieve a near-optimal solution for the entire data set. Coresets are an important tool in scalable machine learning, providing a means to achieve this objective. Coresets have found application in various problem domains, such as clustering (Bundefineddoiu et al., 2002; Har-Peled & Mazumdar, 2004; Frahling & Sohler, 2005, 2008; Feldman & Langberg, 2011; Feldman & Schulman, 2012; Feldman et al., 2013; Braverman et al., 2019; Huang & Vishnoi, 2020; Feldman et al., 2020), linear regression (Drineas et al., 2006; Dasgupta et al., 2009; Clarkson et al., 2019), principal component analysis (Cohen et al., 2015; Feldman et al., 2020), and more (Bachem et al., 2017; Sener & Savarese, 2018; Munteanu et al., 2018; Sohler & Woodruff, 2018; Phillips & Tai, 2018b; Huang et al., 2018; Assadi et al., 2019; Mussay et al., 2020).
Improving the prior result by (Feldman & Langberg, 2011) which suggests that common sensitivity sampling requires $O(dS^{2})$ samples, (Braverman et al., 2021) and (Feldman et al., 2020) proved that $O(dS\log S)$ would be sufficient. Here, $S$ represents the total sensitivity, and $d$ is some VC-dimension associated with the function space. The proof of these results benefits from a connection between approximating a class of functions and approximating their linked-ranges spaces, the same concept established by (Joshi et al., 2011) in the study of $\varepsilon$-coresets for kernel density estimates. It is important to note that these results are subject to the assumption that the space is finite. We eliminate this constraint in Theorem 1.2 with a different and more straightforward proof outlined in Section 2.2. It is notable to highlight that the coreset bounds derived from this theorem are all dependent on the dimension. More importantly, we provide a counterpart of this theorem (see Theorem 1.1) using Rademacher complexity in place of VC-dimension, marking the first result of this kind. As a result, we are able to present several coreset bounds independent of the dimension – the first such results of this type (see Table 2). In the following, we focus more on the related works pertaining to these findings.
Our work shares some connections with the research conducted by (Munteanu et al., 2018; Mai et al., 2021; Tolochinksy et al., 2022), and (Curtin et al., 2020). (Munteanu et al., 2018) demonstrated the existence of datasets for which coresets of sublinear size do not exist. They introduced a complexity measure $\mu(X)$ for data points $X$, quantifying the difficulty associated with compressing a dataset for logistic regression. They demonstrated the existence of coresets with size $O^{*}(\frac{d^{3/2}\mu(X)\sqrt{|X|}}{\varepsilon^{2}})$ and $O^{*}(\frac{d^{3}\mu^{2}(X)}{\varepsilon^{4}})$ through a random sampling procedure $\O^{*}(\cdot)$ hides some logarithmic factors in terms of the problem’s parameters). As a downside of their method, it is not clear how to compute $\mu(X)$ and also they conjectured that computing the value of $\mu(X)$ in general is hard. Moreover, their second bound is roughly quadratic in terms of $\mu(X)$ and $\frac{1}{\varepsilon^{2}}$ when the data possesses a small $\mu$-complexity. Following their work, (Mai et al., 2021) improved these bounds to $O^{*}(\frac{d\mu^{2}(X)}{\varepsilon^{2}})$ which is linear in terms of $\frac{1}{\varepsilon^{2}}$. Their method relies on subsampling data points with probabilities proportional to their $\ell_{1}$ Lewis weights. As they utilize the complexity measure $\mu(X)$ in their context, which cannot be directly interpreted in terms of our variables, presenting a meaningful comparison between our findings and theirs becomes challenging. Nevertheless, as an advantage, some of our bounds are independent of dimension, surpassing their results in this particular aspect.
The most closed works to ours are (Tolochinksy et al., 2022) and (Curtin et al., 2020). With an almost similar setting as ours, but restricted to Euclidean space and with a bounded parameter space, (Tolochinksy et al., 2022) found the coresets of size $O^{*}\left(\frac{kd^{2}\log n\log k}{\varepsilon^{2}}\right)$ for $\phi=\operatorname{{\rm logistic}},\sigma,\operatorname{{\rm svm}}$. Enhancing some of these results, (Curtin et al., 2020) conducted an analysis of sampling-based coreset constructions designed for regularized loss minimization problems (logistic regression or SVM). Their context is slightly distinct from ours. To rephrase their setting in the context of ours, we can express that in our setting, the regularization term is $\frac{1}{k}$ whereas they roughly use $\frac{1}{n^{1-\kappa}}$ for it (up to a constant). In the scenario where $k$ is proportional to $n^{1-\kappa}$ for a fixed $\kappa\in(0,1)$, they have demonstrated that a uniform sample of size $O^{*}\left(\frac{dn^{1-\kappa}}{\epsilon^{2}}\right)$ functions as a coreset with high probability. Here, $n=|X|$ and $d$ represents a type of VC-dimension linked to the function space (refer to the paper for the precise definition).They have further established the tightness of uniform sampling, accurate up to poly-logarithmic factors. A special case of our results not only reproduces their results but also improves upon theirs for logistic regression. It is still noteworthy that their bounds are dimension-dependent, whereas we present some bounds independent of dimension. For a comprehensive comparison, refer to Table 1.
A few other variants of coresets for classification exist; we briefly mention a few. Munteanu et al. (2021) applies sketching to create data compression for logistic regression. Their methods are data oblivious and show improvements for sparse high-dimensional vectors. These results are still polynomial in data parameter $\mu(X)$ and dimension $d$. Mirzasoleiman et al. (2020) considers coresets for regularized classification in the context of incremental gradient descent. They greedily select a coreset to be used in IGD, and their results use that the $\frac{1}{k}\|w\|^{2}$ regularizer makes loss function strongly convex as a function of $k$. Coresets can also be found for the $0/1$ mis-classification function of size $O(d/\varepsilon^{2})$ via classic VC-dimension arguments, but as this cost function is non-convex over the model parameters $w$, the best known algorithms (Matheny & Phillips, 2021) to solve for the approximately optimal solution are still exponential in dimension $O^{*}(1/\varepsilon^{d})$. Or if a dataset is known to be linearly separable, then greedy coresets based on Frank-Wolfe can be found of size roughly $O(1/\varepsilon)$ which approximately preserve that separation margin (Tsang et al., 2005; Gärtner & Jaggi, 2009).
Appendix C Main Tools
@@
C.1 Proof of Theorem 1.1
We begin by revisiting McDiarmid’s inequality, which serves as a concentration inequality that bounds the difference between the sampled mean and the true mean of a specific function. If $(\mathcal{X}_{1},\Sigma_{1},P_{1}),\ldots,(\mathcal{X}_{n},\Sigma_{1},P_{n})$ are probability measure spaces, then $P=\prod_{i=1}^{n}P_{i}$ is a probability measure over $\prod_{i=1}^{n}\mathcal{X}_{i}$. A function $f:\prod_{i=1}^{n}\mathcal{X}_{i}\longrightarrow\mathbb{R}$ satisfies the bounded differences property if there are constants $c_{1},c_{2},\dots,c_{n}$ such that for all $i\in[n]$, and for all $x_{1}\in{\mathcal{X}}_{1},\,x_{2}\in{\mathcal{X}}_{2},\,\ldots,\,x_{n}\in{\mathcal{X}}_{n}$
@@
${\text{P}}\left(f(X_{1},X_{2},\ldots,X_{n})-\mathbb{E}[f(X_{1},X_{2},\ldots,X_{n})]\geq\varepsilon\right)\leq\exp\left(-{\frac{2\varepsilon^{2}}{\sum_{i=1}^{n}c_{i}^{2}}}\right).$
@@
Theorem C.2 (Theorem 1.1, Restated).
@@
Proof.
In view of Equation (3), we need to show that, for any iid sample $x_{1},\ldots,x_{m}\in\mathcal{X}$ according to the sensitivity normalized distribution $q(x)=\frac{s(x)}{S}p(x)$, with probability at least $1-2\exp\left(-{\frac{2mt^{2}}{S^{2}}}\right)$,
@@
((*$\qquad$( ∗ ) $=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{x^{\prime}_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{f\in\mathcal{F}}\left[\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\left(T_{f}(x^{\prime}_{i})-T_{f}(x_{i})\right)\right]$
@@
The equality marked by ()(*)( ∗ ) is true since $\sigma_{i}$ indeed exchanges $x_{i}$ and $x_{i}^{\prime}$ together, and since they are iid samples from $q$ it cannot change the expectation. With the same approach, we would have
@@
which concludes $\max\{A,B\}\leq 2\mathcal{R}^{q}_{m}(\mathcal{T}_{\mathcal{F}})$ completing the proof. ∎
C.2 Contraction Lemma
When dealing with Rademacher complexity, we can leverage the helpful lemma known as Talagrand’s Contraction Lemma (refer to Theorem 4.12 in (Ledoux & Talagrand, 1991)). This lemma, in particular, establishes that $\mathcal{R}_{m}(\phi\circ\mathcal{F})\leq L\mathcal{R}_{m}(\mathcal{F})$, where $\phi\colon\mathbb{R}\longrightarrow\mathbb{R}$ is an $L$-Lipschitz function, and $\phi\circ\mathcal{F}=\{\phi\circ f\colon f\in\mathcal{F}\}$. Our requirement extends beyond this lemma to a generalization.
Lemma C.3.